Pearl Corp. enters into a contract with a customer to build an apartment building for $1,000,800. The customer hopes to rent apartments at the beginning of the school year and provides a performance bonus of $135,600 to be paid if the building is ready for rental beginning August 1, 2018. The bonus is reduced by $45,200 each week that completion is delayed. Pearl commonly includes these completion bonuses in its contracts and, based on prior experience, estimates the following completion outcomes:

Completed by Probability
August 1, 2018 70 %
August 8, 2018 20
August 15, 2018 5
After August 15, 2018 5


Determine the transaction price for this contract.

Transaction Price
Answer #1
Completion by Transaction Amount Probability Expected amount
August 1, 2018 1136400 70% 795480
August 8, 2018 1091200 20% 218240
August 15, 2018 1046000 5% 52300
After August 15, 2018 1000800 5% 50040
Total 1116060
Transaction Price 1116060
Workings:
Completion by Transaction Amount
August 1, 2018 =1000800+135600
August 8, 2018 =1000800+(135600-45200)
August 15, 2018 =1000800+(135600-45200-45200)